# FB-Bench: A Fine-Grained Multi-Task Benchmark for Evaluating LLMs’ Responsiveness to Human Feedback

### v1

@@ 4 Related Work
The evaluation of LLMs is essential for their development. It reveals the strengths and weaknesses of existing models and offers key insights and directions for future research. However, most existing studies [18, 8, 9, 11, 10] focus solely on evaluating the general or specific capabilities of LLMs in single-turn dialogues. They fail to assess LLM performance under various user feedback, which typically involves multi-turn dialogue scenarios. Although there are some benchmarks for multi-turn LLMs [12, 13, 14, 30, 31], the user inputs in the multi-turn dialogues are often independent, lacking feedback towards to the previous LLM output. Furthermore, much of the data in these multi-turn dialogue benchmarks is synthesized by LLMs, failing to exhibit the diversity and complexity of real-world scenarios.
@@
Human feedback not only enhances model performance but also serves as a critical mechanism for aligning the model with desired outcomes or goals [32]. Training models on feedback data, not only can directly enhance the quality of the generated content [33] but also allows models to better align with human preferences in style and tone [34]. During the inference stage, users can provide feedback on intermediate responses, enabling the model to refine its output until it achieves the user’s satisfaction [4, 5]. However, a systematic benchmark for evaluating the impact of human feedback on LLMs during the inference stage is still lacking.
@@
Several benchmarks have begun to explore the impact of feedback on LLMs. However, they predominantly focus on specific tasks or domains. MINT [15] exclusively assesses the coding and reasoning capabilities of LLMs that utilize tools and receive AI-generated language feedback. Intercode [16] evaluates the coding skills of LLMs based on feedback from compilers or interpreters executing the code. AgentBench [17] examines the reasoning and decision-making abilities of LLMs-as-Agents in response to environmental feedback. Different from prior works, FB-Bench introduces a novel approach by measuring the responsiveness of LLMs to diverse user feedback across a broad spectrum of real-world usage scenarios.
@@ 5 Conclusion
We introduce FB-Bench, a fine-grained muti-task benchmark for comprehensively evaluating the responsiveness of LLMs to various human feedback across real-world usage scenarios. A three-tier hierarchical taxonomy, grounded in two popular human-LLM interaction scenarios, is established to ensure thorough coverage of diverse interaction types and scenarios. To facilitate a fine-grained and accurate evaluation, a LLM-as-a-Judge framework, equipped with a weighted checklist, is employed. Benchmarking results from 31 well-known LLMs demonstrate significant performance variations between error correction and response maintenance. Further analysis explores the principal factors influencing the responsiveness of LLMs and provides valuable insights for subsequent research.
@@ A.1.2 Detailed data statistics
The distribution of task and subtask categories is shown in Figure 8.
@@
The length distribution of the four components of conversation in FB-Bench, namely, the user query, the preset model response, user feedback, and the reference follow-up response, is depicted in Figure 9.
Figure 9: The length distribution of four components in conversations within FB-Bench.
@@ A.1.3 Example in FB-Bench
We select an example from error correction and response maintenance scenarios, and display them in Figure 10 and Figure 11, respectively.
Figure 10: The error correction example in FB-Bench.
Figure 11: The response maintenance example in FB-Bench.
A.2 Detailed description of experiments
A.2.1 Response generation
We utilize the vllm library [35] to deploy open-source LLMs for generating follow-up responses based on a user query, a predetermined model response, and human feedback. In terms of temperature settings, we assign distinct values for different tasks: 0.7 for text creation and text translation, 0.1 for knowledge-based Q&A, and 0 for all other tasks. For the maximum output length, we set it to the minimum value between 4096 and the difference between the LLM context length and the context tokens length.
A.2.2 Evaluation
@@
A.2.3 The full results in FB-Bench
We evaluate 31 popular LLMs using FB-Bench, with the results presented in Table 3.
@@ A.2.3 The full results in FB-Bench
Model Error Correction Response Maintenance Overall
qwen-max-0919 71.52 62.57 67.04
Qwen2.5-72B-Instruct 66.51 63.22 64.87
ERNIE-4.0-8K-0329 66.30 62.59 64.44
gpt-4o-2024-05-13 69.90 55.01 62.46
gpt-4-turbo-2024-04-09 67.24 56.08 61.66
Baichuan-Instruct 65.65 57.30 61.48
glm-4-0520 66.40 55.30 60.85
Qwen2-72B-Instruct 63.46 57.81 60.63
claude-3-5-sonnet-20240620 73.87 46.34 60.11
gpt-4o-mini-2024-07-18 66.74 50.55 58.65
internlm2_5-20b-chat 55.47 60.30 57.89
yi-large 63.28 50.91 57.10
Mistral-Large-Instruct-2407 64.69 46.30 55.50
moonshot-v1-32k 59.57 51.41 55.49
DeepSeek-V2.5 64.47 46.35 55.41
Qwen2.5-7B-Instruct 55.75 49.00 52.37
internlm2_5-7b-chat 49.66 54.37 52.01
Qwen2-7B-Instruct 48.05 50.71 49.38
Yi-1.5-34B-Chat-16K 48.17 46.21 47.19
Yi-1.5-9B-Chat-16K 45.65 47.95 46.80
claude-3-sonnet-20240229 60.48 31.98 46.23
glm-4-9b-chat 58.60 33.49 46.04
Yi-1.5-34B-Chat 50.55 40.69 45.62
DeepSeek-Coder-V2-Lite-Instruct 53.40 37.10 45.25
Phi-3-medium-4k-instruct 36.17 54.29 45.23
Meta-Llama-3.1-70B-Instruct 56.05 32.26 44.16
gemma-2-27b-it 54.31 31.69 43.00
Yi-1.5-9B-Chat 47.91 34.42 41.16
Phi-3-small-8k-instruct 30.98 47.87 39.43
gemma-2-9b-it 51.58 24.95 38.27
Meta-Llama-3.1-8B-Instruct 47.31 26.26 36.78

### v2

@@ 4 Related Work
The evaluation of LLMs is essential for their development. It reveals the strengths and weaknesses of existing models and offers key insights and directions for future research. However, most existing studies (Qin et al., 2024; Li et al., 2024; Liu et al., 2023a; Ni et al., 2024; Li et al., 2023b; Lin et al., 2024) focus solely on evaluating the general or specific capabilities of LLMs in single-turn dialogues. They fail to assess LLM performance under various user feedback, which typically involves multi-turn dialogue scenarios. Although there are some benchmarks for multi-turn LLMs (Zheng et al., 2024; Sun et al., 2024; Bai et al., 2024; Kwan et al., 2024; Liang et al., 2024), the user inputs in the multi-turn dialogues are often independent, lacking feedback towards to the previous LLM output. Furthermore, much of the data in these multi-turn dialogue benchmarks is synthesized by LLMs, failing to exhibit the diversity and complexity of real-world scenarios.
@@
Human feedback not only enhances model performance but also serves as a critical mechanism for aligning the model with desired outcomes or goals (Wiener, 2019). Training models on feedback data, not only can directly enhance the quality of the generated content (Ouyang et al., 2022) but also allows models to better align with human preferences in style and tone (Ziegler et al., 2019). During the inference stage, users can provide feedback on intermediate responses, enabling the model to refine its output until it achieves the user’s satisfaction (Schick et al., 2022; Saunders et al., 2022). However, a systematic benchmark for evaluating the impact of human feedback on LLMs during the inference stage is still lacking.
@@
Several benchmarks have begun to explore the impact of feedback on LLMs. However, they predominantly focus on specific tasks or domains. MINT (Wang et al., 2023) exclusively assesses the coding and reasoning capabilities of LLMs that utilize tools and receive AI-generated language feedback. Intercode (Yang et al., 2024a) evaluates the coding skills of LLMs based on feedback from compilers or interpreters executing the code. AgentBench (Liu et al., 2023b) examines the reasoning and decision-making abilities of LLMs-as-Agents in response to environmental feedback. Different from prior works, FB-Bench introduces a novel approach by measuring the responsiveness of LLMs to diverse user feedback across a broad spectrum of real-world usage scenarios.
@@ 5 Conclusion
We introduce FB-Bench, a fine-grained muti-task benchmark for comprehensively evaluating the responsiveness of LLMs to various human feedback across real-world usage scenarios in Chinese. A three-tier hierarchical taxonomy, grounded in two human-LLM interaction scenarios, is established to ensure thorough coverage of diverse interaction types and scenarios. To facilitate a fine-grained and accurate evaluation, a LLM-as-a-Judge framework, equipped with a weighted checklist, is employed. Benchmarking results from 27 well-known LLMs demonstrate significant performance variations between error correction and response maintenance. Further analysis explores the principal factors influencing the responsiveness of LLMs and provides valuable insights for subsequent research.
Limitations
Here we discuss several limitations of this work.
  • Although our data curation pipeline and evaluation framework are designed to be flexible and adaptable for use across various languages, the FB-Bench dataset is currently available only in Chinese. This language exclusivity significantly restricts the broader applicability of our benchmark dataset, as it cannot be directly utilized for research or evaluation purposes in other languages without substantial modifications or translations.
  • Our evaluation protocol primarily relies on LLMs. Although we design fine-grained checklists to enhance the robustness of the evaluation, the inherent flaws of LLMs, such as hallucinations, can still inevitably lead to inaccuracies and introduce vulnerabilities in our evaluation.
@@ A.1.2 Detailed data statistics
The distribution of task and subtask categories in our dataset is shown in Figure 8.
@@
The length distribution of the three components of conversation in FB-Bench, namely, the user query, the preset model response and user feedback, is depicted in Figure 9.
Figure 9: The length distribution of three components in conversations within FB-Bench.
@@ A.1.3 Example in FB-Bench
We select examples from the error correction and response maintenance scenarios, which are displayed in Figure 10.
Figure 10: Examples wthin error correction and response maintenance scenarios in FB-Bench.
A.2 Detailed description of data curation
A.2.1 Heuristic rules
For error correction data, we initially filter target data from online chat data using the following heuristic rules:
  • Multi-turn dialogues
  • Created within the past three months
  • The model’s second-turn response contains phrases such as “I’m sorry”, “I apologize”, or “Please forgive me”.
  • Token length is less than 5000
  • Does not use retrieval augmentation
  • No security restrictions are triggered
Subsequently, we include samples with scores below 3 from human preference data, where the model’s responses are deemed poor in quality. Scores for human preference data range from 1 to 5, with higher scores indicating better quality.
For response maintenance data, we include answers and corresponding dialogues that have received user upvotes by analyzing the front-end log records. Additionally, we remain data with scores above 4 from the human preference dataset.
A.2.2 Data annotation
We employ a two-stage labeling process, which includes GPT pre-annotation followed by human annotation, to label each sample in the collected dataset.
GPT pre-annotation
For each sample within error correction scenario, GPT-4o-2024-05-13(denoted as GPT-4o for convenience in the following text) is utilized to generate a series of preliminary labels. This process involves using GPT-4o to produce a reference response to the user’s query. Subsequently, GPT-4o identifies the reason for user dissatisfaction and simulates a user providing detailed feedback, assigning a corresponding predefined feedback category. Following this, GPT-4o generates a follow-up response based on the user query, the model’s response, and the feedback provided. Finally, GPT-4o generates an instance-level checklist consisting of several yes/no questions, each associated with a weight ranging from 0 to 1, where a higher weight signifies greater importance. The cumulative weight of these questions equals 1. For each sample within response maintenance scenario, since the model response is good enough, we employ GPT-4o to simulate a real user providing several detailed feedback for each of the predefined feedback types. We then ask GPT-4o to generate an instance-level checklist, which also consists of several yes/no questions, albeit without associated weights.
Human annotation
To further improve the diversity and quality of the dataset, we engage ten Chinese annotators, each holding a bachelor’s degree, to review, revise, and refine the data initially annotated by GPT-4o. To provide guidance on data annotation, we initially prepare a document detailing the dataset’s background, the meaning of each field, the value ranges, and specific examples. Subsequently, we organize the data into a table and annotate three entries for each task. Finally, we conduct a conference to instruct human annotators on data annotation, integrating the document and sample data. Following the guidance, human annotators are asked to randomly select and annotate five entries per task. We then review these annotations to ensure quality and consistency, confirming their understanding of our requirements. In the labeling process for each sample in the error correction scenario, human annotators first assess whether the task, subtask, and deficiency types are misaligned. They then refine or rewrite the human feedback and its corresponding feedback type, followed by refining or rewriting the associated checklist. This process involves modifying all content related to yes/no questions and the corresponding weights. For each sample in the response maintenance scenario, annotators similarly begin by checking the task and subtask types. They then provide the human feedback and its corresponding feedback type, concluding with the checklist, as in the error correction scenario. Each sample in FB-Bench is annotated by one annotator and subsequently reviewed by another.
A.3 Detailed description of experiments
A.3.1 Response generation
We utilize the vllm library (Kwon et al., 2023) to deploy open-source LLMs for generating follow-up responses based on a user query, a preset model response, and human feedback. In terms of temperature settings, we assign distinct values for different tasks: 0.7 for text creation and text translation, 0.1 for knowledge-based Q&A, and 0 for all other tasks. For the maximum output length, we set it to the minimum value between 4096 and the difference between the LLM context length and the context tokens length.
A.3.2 Evaluation
@@
A.4 Agreement Evaluation
To justify the weighted checklist LLM-as-a-Judge method we designed for FB-Bench, we conduct extensive human evaluation over FB-Bench’s randomly selected examples.
We collect 194 samples by performing random sampling for each task within FB-Bench, along with the following-up responses from five LLMs: claude-3-5-sonnet-20240620, gpt-4o-2024-05-13, qwen-max-0919, Qwen2.5-72B-Instruct, and internlm2_5-20b-chat. We then replace the selected models’ names with Model1 to Model5 and ask human annotators to evaluate the following-up responses based on the corresponding checklist. Subsequently, we employ several advanced LLMs, including including gpt-4o-2024-08-06, gpt-4o-2024-11-20, qwen-max-0919, and claude-3-5-sonnet-20241022, as judge models to perform evaluations as well. Finally, we compile all the judgment results generated by human annotators and the judge models to calculate the consistency rate.
Table 4 presents the consistency rate between human annotators and several judge models. The consistency rate for gpt-4o-2024-08-06 is the highest, at 84.79%. Consequently, we select gpt-4o-2024-08-06 as the judge model in FB-Bench. The corresponding score of the five selected LLMs under different judge models can be found in Table 5.
A.4.1 The full results in FB-Bench
We evaluate 27 popular LLMs using FB-Bench, with the results presented in Table 3.
Model Error Correction Response Maintenance Overall
DeepSeek-V3 74.24 75.48 74.86
qwen-max-0919 72.03 74.42 73.22
gpt-4o-2024-11-20 71.74 73.68 72.71
Qwen2.5-72B-Instruct 67.95 71.29 69.62
ERNIE-4.0-8K-0329 62.37 75.85 69.11
claude-3-5-sonnet-20241022 69.92 67.31 68.61
Mistral-Large-Instruct-2411 64.97 70.36 67.66
yi-lightning 70.09 64.69 67.39
glm-4-0520 60.25 73.43 66.84
gpt-4o-2024-05-13 68.82 64.72 66.77
claude-3-5-sonnet-20240620 68.99 63.95 66.47
phi-4 58.17 74.67 66.42
internlm3-8b-instruct 50.85 81.04 65.95
moonshot-v1-32k 59.09 69.27 64.18
gpt-4o-mini-2024-07-18 61.83 63.74 62.79
internlm2_5-20b-chat 53.44 71.79 62.61
Qwen2.5-7B-Instruct 57.09 58.62 57.85
internlm2_5-7b-chat 48.14 67.45 57.79
glm-4-9b-chat 58.60 54.92 56.76
Yi-1.5-34B-Chat 46.98 60.26 53.62
Meta-Llama-3.1-70B-Instruct 54.51 52.69 53.60
Phi-3-medium-4k-instruct 37.62 65.54 51.58
Ministral-8B-Instruct-2410 49.77 51.55 50.66
gemma-2-27b-it 48.55 51.24 49.89
Yi-1.5-9B-Chat 43.62 54.39 49.00
gemma-2-9b-it 44.56 46.84 45.70
Meta-Llama-3.1-8B-Instruct 51.17 38.21 44.69
@@ A.4.1 The full results in FB-Bench
Judge Model Consistency Rate
gpt-4o-2024-08-06 90.91%
gpt-4o-2024-11-20 88.77%
qwen-max-0919 89.42%
claude-3-5-sonnet-20241022 90.47%
Table 4: The consistency rate between human and different judge models.
gpt-4o-2024-05-13 claude-3-5-sonnet-20240620 internlm2_5-20b-chat Qwen2.5-72B-Instruct qwen-max-0919
gpt-4o-2024-08-06 75.11 73.88 72.28 70.71 66.59
gpt-4o-2024-11-20 74.61 75.79 71.99 69.99 60.62
qwen-max-0919 77.15 77.76 73.04 75.05 66.36
claude-3-5-sonnet-20241022 73.96 76.07 71.02 72.59 62.42
Table 5: The overall score of the five selected LLMs under different judge models. The header represents the evaluated models, while the index indicates the judges.